# LLMs generate structurally realistic social networks but overestimate political homophily

### v1

@@ Appendix A Comparison between LLMs
(a) Comparing structural network metrics.
(b) Comparing homophily.
Figure 7: Evaluating results from GPT-4o, compared to GPT-3.5 Turbo.
(a) Comparing structural network metrics.
(b) Comparing homophily.
Figure 8: Evaluating results from Llama 3.1 8B and 70B, compared to GPT-3.5 Turbo.
(a) Comparing structural network metrics.
(b) Comparing homophily.
Figure 9: Evaluating results from Gemma 2 9B and 27B, compared to GPT-3.5 Turbo.
We experiment with the following LLMs: OpenAI’s GPT-3.5 Turbo and GPT-4o Brown et al. (2020); OpenAI et al. (2023), Meta’s Llama 3.1 (8B and 70B) Touvron et al. (2023), and Google’s Gemma 2 (9B and 27B) Gemma Team et al. (2024), representing a range of models across companies and sizes. For the OpenAI models, we use the OpenAI API, [^5] https://platform.openai.com/docs/api-reference and for the other models, we use the Llama API, [^6] https://www.llama-api.com/ which also includes other open-source models. We report our main results on GPT-3.5 Turbo since we find that it best matches the structure of the real-world social networks (discussed below), but here, we discuss results from all other models. For these experiments, we generate networks with the Sequential method, using the same experimental settings as before (same set of 50 personas, prompt, temperature of 0.8, etc.). For these experiments, we generate 10 instead of 30 networks per model, but we find that standard errors are small. We visualize results for GPT-4o in Figure 7, Llama 3.1 8B and 70B in Figure 8, and Gemma 2 9B and 27B in Figure 9.
@@
We find that GPT-3.5 Turbo best matches the structure of the real networks, most notably matching the real-world density. All of the other models have much higher densities, which also contributes to unrealistically high clustering and low shortest paths. Due to density’s outsized effect, we wanted to see if providing the model a bit of help on density might be all that it needs to match the other characteristics as well. Thus, we try a variant of the Sequential method where we specify $n$, i.e., how many friends should be chosen (instead of only asking “Which of these people will you become friends with?”, see full prompt in Figure 19). We sample $n$ from Exponential($\lambda=0.2$), with mean $51/\lambda=$ 5, independently for each persona. Note that specifying each persona’s number of choices does not predetermine the exact density or degree distribution of the network, since a persona’s total set of friends at the end of the network generating process is the union of its chosen friends along with anyone who chose it. However, specifying these numbers can help to guide the models to lower densities.
With this variant, which we call “$+\lambda$”, all models’ densities are brought down to a reasonable range. Llama 3.1 8B $+\lambda$ is also able to approximately match the real networks and GPT-3.5 Turbo on all other characteristics now (Figure 8). However, even with $+\lambda$, GPT-4o and Llama 3.1 70B still generate networks that are often disconnected into two components; this is due to extreme political segregation, which we discuss below. With Gemma 2 27B, the networks are almost always fully connected, but even with $+\lambda$, clustering and modularity remain slightly too high. [^7] We were not able to add $+\lambda$ to Gemma 2 9B, since it had trouble producing responses that exactly matched $n$.
@@
We find that all six models exhibit clear homophily, with ratios above 1 for gender, race/ethinicity, religion, and political affiliation. We also find, consistently across the models, that political homophily remains by far the strongest, followed by religion, then race/ethnicity and gender. Political homophily is also always overestimated compared to real-world measures; in fact, the models we test in this section all produce levels of political homophily even higher than GPT-3.5 Turbo, which we discussed were overestimates in the main text (Table 1).
The emphasis on political homophily is particularly extreme for GPT-4o and Llama 3.1 70B: they show complete segregation between Republicans and Democrats, such that 100% of edges are same-group and political homophily reaches its maximum value. We can compute the maximum value based on the known numbers of nodes in each group: in our set of 50 sampled personas, we have 26 Republicans and 24 Democrats (Table 7). So, the expected proportion of same-group edges is $\frac{(26\cdot 25)+(24\cdot 23)}{50\cdot 49}\approx 0.49$ (Eq. 1), so the maximum ratio is $1/0.49\approx 2.04$, which is indeed the ratio that we see for GPT-4o and Llama 3.1 70B (with and without $+\lambda$). Furthermore, as a result of the complete segregation, their networks fracture into two disconnected components, one for Republicans and one for Democrats, which is why we see that the proportion of nodes in the largest connected component is often around 50% for these two models (Figure 7 and 8).
@@ Appendix B Additional results
Table 3: Quantitative results on structural characteristics. The top six rows indicate difference in means, normalized by the real networks’ standard deviation (Eq. 4). indicates that the model parameters were fitted on the real-world mean value for this characteristic, so the mean difference should be ignored, and it is left out of the average. The bottom six rows represent the two-sample Kolmogorov–Smirnov statistic, which measures the distance between two empirical distributions (Eq. 5). For both measures, lower is better.
@@
In the main text, we showed that the LLM can match many structural characteristics of real social networks, including density, clustering, connectivity, and degree distribution (Figures 3 and 4). However, how does this compare to existing models for network generation? Here, we consider three classical network models: (1) Erdős–Rényi random graph models Erdös and Rényi (1959), (2) Barabási-Albert preferential attachment models Barabási and Albert (1999), and (3) Watts-Strogatz small-world models Watts and Strogatz (1998). These models are close to zero-shot: they only require a 1-2 parameters to define, although we still need to fit these parameters. In contrast, we do not compare to stochastic block models Holland et al. (1983), which require sizes per block and probabilities between each pair of blocks, or machine learning models for graph generation, since they have many parameters and require a substantial set of observed graphs for training to fit those parameters You et al. (2018); Simonovsky and Komodakis (2018); Guo and Zhao (2023).
@@
To quantify how well a network metric is matched, we extract the metric from each real social network, using the same eight social networks as in our main experiments (Sections 5.1 and C.3), and from each generated network, with 30 generated networks per model. To compare the two distributions of the metric, we report both the difference in the means, normalized by the real networks’ standard deviation (Eq. 4), and the two-sample Kolmogorov–Smirnov statistic (Eq. 5), which measures the distance between two empirical distributions. As we show in Table 3, even in the best case—when we allow the models to choose parameters based on the real social networks—these models cannot match all of the real network metrics as well as our Local or Sequential methods can. Thus, being able to match the structural characteristics of real social networks is non-trivial, adding significance to our finding that LLMs can match many structural characteristics at once.
@@
Figure 10: Observed (left) vs. expected (right) distribution of age gaps in the generated social networks, under the Sequential method.
We cannot directly apply our typical homophily measure, Eq. 1, to age, since it is not a categorical variable. However, we can still measure age homophily in a number of ways. First, we can replace $\mathbbm{1}[g_{i}=g_{j}]$ with $\mathbbm{1}[|age(i)-age(j)|\leq 10]$ in Eq. 1. Then, the ratio we are computing is the proportion of observed edges with an age gap less than or equal to 10, divided by the expected proportion of such edges. Note that ages range from 0 to 100 (Section C.1), so a gap of 10 is relatively small. Under this measure, we find values of 1.376 for Global, 1.254 for Local, and 1.258 for Sequential, indicating clear homophily. We can also plot the distribution of observed age gaps vs. expected age gaps, which we show in Figure 10. Compared to the expected distribution, the observed distribution is clearly shifted to the left (i.e., smaller gaps), with more edges than expected with age gaps of 0-10.
Reasons for choosing friends.
Demographic Reason %
Political affiliation 86.7%
Religion 43.0%
Age 21.8%
Race/ethnicity 12.1%
Gender 7.3%
Table 4: Frequency that each demographic is part of the reason given for choosing a persona as a friend.
We have seen under different models and prompts that political homophily is the strongest, out of the demographic variables that we test. We have also shown, after shuffling demographics to remove correlations between them, that the LLM still favors political homophily the most (Figure 5), suggesting that it is, in fact, paying the most attention to political affiliation when forming social relations. As another way to test this theory, we try directly asking the LLM to generate a short reason for each friend it selects. Then, given the reason, we use GPT-4o to classify the reason, e.g., “I’m a woman too and we share the same religion and political affiliation” is classified as [gender, religion, political affiliation]. We allow the LLM to generate free-text reasons during network generation since we do not want to constrain its response (e.g., if it is using other information, such as degree or ID).
We find that political affiliation strongly dominates here as well: it is part of the reason 86.7% of the time, while the next most-mentioned demographic, religion, is only mentioned 43.0% of the time (Table 4). As a caveat, note that we cannot always trust an LLM’s self-given reason or explanation for its choices Agarwal et al. (2024). However, given the alignment of these results along with everything else we have seen, there is strong evidence that LLMs pay the most attention to political affiliation when generating social relations. We also test whether prompting the LLM to give a reason changes the structure of the generated networks. We show in Figure 14 that our main results do not significantly change, so they are robust to this modification in prompt.
Homophily between all pairs of groups.
Figure 11: Ratio of observed-to-expected proportion of edges, for all pairs of demographic groups (Eq. 2). All subfigures share the same colormap (right). Groups with at least 3 nodes are kept.
As described in the main text, we measure homophily as the ratio of observed same-group relations to expected same-group relations, based on the number of nodes in each group (Eq. 1). Now, we extend this definition to compute a measure for any pair of groups, $A$ and $B$ (where $A$ could equal $B$):
$H_{AB}$ $=\frac{\frac{\sum_{ij}A_{ij}\cdot\mathbbm{1}[g_{i}=A]\cdot\mathbbm{1}[g_{j}=B]}{E}}{\frac{N_{A}(N_{B}-\mathbbm{1}[A=B])}{N(N-1)}}.$ (2)
This measures the ratio of observed to expected edges for pair $A,B$. In Figure 11, we visualize $H_{AB}$ for all pairs, for each of the four categorical demographic variables (gender, race/ethnicity, religion, and political affiliation). For each variable, we keep all groups with at least 3 personas in our set of 50 personas (Table 7), and compute $H_{AB}$ for each pair of groups, reporting the mean ratio over the 30 generated networks from Sequential.
We find, as expected, that the diagonal (i.e., same-group relations) tends to be above 1, although notably it is not for relations between men, with a ratio of 0.99, while relations between women have a ratio of 1.43. Thus, even within one demographic variable, the LLM’s levels of homophily vary for different groups, such as homophily within women being stronger than homophily within men. Variability across groups is present for all demographic variables: most extremely, for race/ethnicity, relations between personas that identify as Black occur 3.44$\times$ more frequently than expected, the highest ratio we find, while overall homophily for race/ethnicity (following Eq. 1) is only 1.22. For political affiliation, both parties have strong same-group tendencies, with ratios of 1.54 and 1.85 for Republicans and Democrats, respectively, and a low cross-group ratio of 0.34.
@@
Figure 12: Visualizing demographic homophily under three methods: “Sequential” (only demographic information), “Sequential, w interests” (demographic information and interests), and “Sequential, only interests” (only interests). The LLM places the largest emphasis on political homophily in all cases.
@@
Table 5: Top 10 interests per demographic group.
@@
In Figure 12, we visualize demographic homophily under three versions of the Sequential method: when only demographic information is provided, when demographic information and interests are provided, and when only interests are provided. In all cases, the LLM places the largest emphasis on political homophily.

### v2

@@ Appendix A Comparison Between LLMs
Figure A1: Evaluating results from GPT-4o, compared to GPT-3.5 Turbo. Top: structural network metrics. Bottom: homophily, where ratios below 1 (marked by the grey line) indicate homophily and lower ratios indicate more homophily. In both plots, we visualize mean and standard error (in black) and individual data points corresponding to each network.
Figure A2: Evaluating results from Llama 3.1 8B and 70B, compared to GPT-3.5 Turbo. Top: structural network metrics. Bottom: homophily, where ratios below 1 (marked by the grey line) indicate homophily and lower ratios indicate more homophily. In both plots, we visualize mean and standard error (in black) and individual data points corresponding to each network.
Figure A3: Evaluating results from Gemma 2 9B and 27B, compared to GPT-3.5 Turbo. Top: structural network metrics. Bottom: homophily, where ratios below 1 (marked by the grey line) indicate homophily and lower ratios indicate more homophily. In both plots, we visualize mean and standard error (in black) and individual data points corresponding to each network.
We experiment with the following LLMs: OpenAI’s GPT-3.5 Turbo and GPT-4o (Brown et al. 2020; OpenAI et al. 2023), Meta’s Llama 3.1 (8B and 70B) (Touvron et al. 2023), and Google’s Gemma 2 (9B and 27B) (Gemma Team et al. 2024). These six models represent a range across companies, different model sizes, and proprietary (the GPT models) versus open-source (the Llama and Gemma models). Llama 3.1 in particular includes some of the best-performing open-source models on holistic benchmarks. [^2] https://crfm.stanford.edu/helm/lite/latest/#/leaderboard. For the OpenAI models, we use the OpenAI API, [^3] https://platform.openai.com/docs/api-reference and for the other models, we use the Llama API, [^4] https://www.llama-api.com/ which also includes other open-source models. We report our main results on GPT-3.5 Turbo since we find that it best matches the structure of the real-world social networks (discussed below), but here, we discuss results from all other models. For these experiments, we generate networks with the Sequential method, using the same experimental settings as before (same set of 50 personas, prompt, temperature of 0.8, etc.). For these experiments, we generate 10 instead of 30 networks per model, but we find that standard errors are small. We visualize results for GPT-4o in Figure A1, Llama 3.1 8B and 70B in Figure A2, and Gemma 2 9B and 27B in Figure A3.
@@
We find that GPT-3.5 Turbo best matches the structure of the real networks, most notably matching the real-world density. All of the other models have much higher densities, which also contributes to unrealistically high clustering and low shortest paths. Due to density’s outsized effect, we wanted to see if providing the model a bit of help on density might be all that it needs to match the other characteristics as well. Thus, we try a variant of the Sequential method where we specify $n$, i.e., how many friends should be chosen (instead of only asking “Which of these people will you become friends with?”, see full prompt in Figure C5). We sample $n$ from Exponential($\lambda=0.2$), with mean $51/\lambda=$ 5, independently for each persona. Note that specifying each persona’s number of choices does not predetermine the exact density or degree distribution of the network, since a persona’s total set of friends at the end of the network generating process is the union of its chosen friends along with anyone who chose it. However, specifying these numbers can help to guide the models to lower densities.
With this variant, which we call “$+\lambda$”, all models’ densities are brought down to a reasonable range. Llama 3.1 8B $+\lambda$ is also able to approximately match the real networks and GPT-3.5 Turbo on all other characteristics now (Figure A2). However, even with $+\lambda$, GPT-4o and Llama 3.1 70B still generate networks that are often disconnected into two components; this is due to extreme political segregation, which we discuss below. With Gemma 2 27B, the networks are almost always fully connected, but even with $+\lambda$, clustering and modularity remain slightly too high. [^5] We were not able to add $+\lambda$ to Gemma 2 9B, since it could not consistently follow instructions to choose exactly $n$ friends.
@@
We find that all six models exhibit clear homophily, with ratios above 1 for gender, age, race/ethnicity, religion, and political affiliation. We also find, consistently across the models, that political homophily remains by far the strongest form of homophily and it is always overestimated compared to real-world measures. In fact, the models we test in this section all produce levels of political homophily even higher than GPT-3.5 Turbo, which we discussed were overestimates in the main text (Table 2).
The emphasis on political homophily is particularly extreme for GPT-4o and Llama 3.1 70B: they show complete segregation between Republicans and Democrats, such that there are no cross-group edges and the cross-group ratio is around 0 (Figures A1 and A2). Furthermore, as a result of the complete segregation, their networks fracture into two disconnected components, one for Republicans and one for Democrats, which is why we see that the proportion of nodes in the largest connected component is often around 50% for these two models. This is highly unrealistic, as it is well-known that social networks, despite having strong community structure and homophily, are also characterized by having a giant connected component that contains the vast majority of the nodes (Ugander et al. 2011). Splitting a synthetic network into two disconnected components has important implications for downstream modeling and use cases of these generated networks: for example, an epidemic outbreak started in one component would never reach the other component, or opinions could not spread, greatly altering the trajectory of dynamic network processes.
@@ Appendix B Additional Results
Table B1: Quantitative results on structural characteristics. The top six rows indicate difference in means, normalized by the real networks’ standard deviation (Eq. 6). indicates that the model parameters were fitted on the real-world mean value for this characteristic, so the mean difference should be ignored, and it is left out of the average. The bottom six rows represent the two-sample Kolmogorov–Smirnov statistic, which measures the distance between two empirical distributions (Eq. 7). For both measures, lower is better.
Study / LLM Method Demographic Cross-group ratio (Eq. 1) $\downarrow$
Thelwall (2009), MySpace Gender 1.04
Race/ethnicity 0.44
Religion 0.69
Laniado et al. (2016), Tuenti, friendship Gender 0.96
Laniado et al. (2016), Tuenti, interaction Gender 0.88
Smith, McPherson, and Smith-Lovin (2014), GSS in 1985 Gender 0.81
Age 0.59
Race/ethnicity 0.17
Religion 0.45
Smith, McPherson, and Smith-Lovin (2014), GSS in 2004 Gender 0.88
Age 0.60
Race/ethnicity 0.25
Religion 0.44
Local Gender 0.774 (0.015)
Age 0.894 (0.010)
Race/ethnicity 0.796 (0.013)
Religion 0.755 (0.013)
Sequential Gender 0.779 (0.017)
Age 0.900 (0.013)
Race/ethnicity 0.804 (0.018)
Religion 0.810 (0.016)
Table B2: Comparing levels of homophily in real-world vs. LLM-generated networks, for the demographic variables besides political affiliation. For Local and Sequential, we report the mean and standard error (in parentheses) over each method’s 30 generated networks.
Comparing to classical network models.
@@
In the main text, we showed that the LLM can match many structural characteristics of real social networks, including density, clustering, connectivity, and degree distribution (Figures 3 and 4). However, how does this compare to existing models for network generation? Here, we consider three classical network models: (1) Erdős–Rényi random graph models (Erdős and Rényi 1959), (2) Barabási-Albert preferential attachment models (Barabási and Albert 1999), and (3) Watts-Strogatz small-world models (Watts and Strogatz 1998). We choose these models to compare to since they only have 1-2 parameters, and we are interested in LLM’s capabilities to generate networks without additional training and with minimal need to defined parameters. In contrast, we do not compare to stochastic block models (Holland, Laskey, and Leinhardt 1983), which require edge probabilities between all pairs of blocks and block assignments, or machine learning models for graph generation, since they have many parameters and require a substantial set of observed graphs for training to fit those parameters (You et al. 2018; Simonovsky and Komodakis 2018; Guo and Zhao 2023).
@@
To quantify how well a network metric is matched, we extract the metric from each real social network, using the same eight social networks as in our main experiments (Sections 5.1 and C.3), and from each generated network, with 30 generated networks per model. To compare the two distributions of the metric, we report both the difference in the means, normalized by the real networks’ standard deviation (Eq. 6), and the two-sample Kolmogorov–Smirnov statistic (Eq. 7), which measures the distance between two empirical distributions. As we show in Table B1, even in the best case—when we allow the models to choose parameters based on the real social networks—these models cannot match all of the real network metrics as well as our Local or Sequential methods can. Thus, being able to match the structural characteristics of real social networks is non-trivial, adding significance to our finding that LLMs can match many structural characteristics at once.
Real-world homophily for other demographics.
In the main text, we showed that levels of political homophily predicted by the LLM are unrealistically high, compared to reported levels of homophily in real-world social networks (Table 2). What about real-world levels of homophily for the other demographic variables? Data from Thelwall (2009), who studies connections on MySpace, suggests cross-group ratios of 0.44 for race/ethnicity, 0.69 for religion, and 1.04 for gender. Data from Laniado et al. (2016), who study connections on Tuenti (a Spanish social network platform), suggests cross-gender ratios of 0.96 when measuring friendship networks and 0.88 when measuring interaction networks. Smith, McPherson, and Smith-Lovin (2014), using data from the US General Social Survey (GSS), finds cross-group ratios of 0.17 for race/ethnicity, 0.45 for religion, 0.59 for age, and 0.81 for gender in 1985 and ratios of 0.25 for race/ethnicity, 0.44 for religion, 0.60 for age, and 0.88 for gender in 2004.
These numbers, which we summarize in Table B2, reveal how much variation there is in levels of homophily across studies and over time, making it difficult to evaluate whether the LLM’s levels of homophily are realistic or not. However, some trends emerge: gender homophily tends to be weaker, race/ethnicity homophily tends to be stronger, and religion and age are somewhere in between. Quotes from summary papers on homophily support this ordering: McPherson, Smith-Lovin, and Cook (2001) say, “By the time that they are adults, people have friendship and confidant networks that are relatively sex-integrated (at least when compared to other dimensions like race, age, and education),” and Thelwall (2009) says, “For U.S. friendship in the last century, the key factors, in decreasing order, seem to be race and ethnicity, age, religion, educational level, occupation, and gender.”
In contrast, in our LLM experiments, the Sequential method predicts similar levels of homophily for gender, race/ethnicity, and religion—all around 0.8—and less homophily for age (0.9) (Table B2). The Local method predicts slightly more homophily for all variables, but a similar ranking: gender, race/ethnicity, and religion have cross-group ratios around 0.75 and age still has less homophily (0.89). While these levels of homophily seem realistic for gender, it seems that the LLM seriously underestimates homophily for race/ethnicity and religion (and likely for age as well, although this is inconclusive since we only have age homophily from one study and age homophily is less comparable to the others since it is not a categorical variable). It is possible that LLMs, due to their guardrails and instruction fine-tuning, have been guided away from using sensitive attributes like race or religion to influence their generation; however, future work is needed to thoroughly investigate evidence of underestimation and probe for possible reasons.
@@
Figure B1: Observed (left) vs. expected (right) distribution of age gaps in the generated social networks, under the Sequential method.
Age is our only non-categorical demographic variable, which slightly complicates our definition of homophily, which assumes well-defined groups. We are able to measure an approximately analogous cross-group ratio (Eq. 1) for age by comparing the observed-to-expected average age gaps between connected nodes, as described in the main text. As a more detailed measure, we can also plot the distribution of observed age gaps vs. expected age gaps, which we show in Figure B1. The observed distribution consists of the age gaps for all observed edges pooled over the 30 networks generated by the Sequential method. The expected distribution consists of the age gaps between all possible pairs of nodes among the 50 personas. Compared to the expected distribution, the observed distribution is clearly shifted to the left (i.e., smaller gaps), with more edges than expected with age gaps of 0-10.
Ablations of demographic variables.
Figure B2: Evaluating homophily per demographic variable, when providing subsets of the demographics. Top: providing one variable at a time. Bottom: providing two variables at a time, political affiliation (PA) and one of the other four. Ratios below 1 (marked by the grey line) indicate homophily and lower ratios indicate more homophily. In both plots, we visualize mean and standard error (in black) and individual data points corresponding to each network. All model results shown here use the Sequential method and GPT-3.5-Turbo.
Here, we explore how the LLM’s generated networks change when we provide subsets of the original set of five demographic variables (gender, age, race/ethnicity, religion, and political affiliation). First, we try providing one variable at a time. When only one variable is provided, we find that homophily for that variable gets stronger, compared to when all five variables are provided (Figure B2, top). This is not entirely surprising: when the LLM is only provided this variable, then it forms the social network based entirely on that variable. For example, if the LLM is only given gender, the LLM is unlikely to connect a woman and a man since they do not have any listed demographics in common, but if more demographics were provided, they might now have a demographic in common. However, this result is not guaranteed: the LLM could have learned the joint distribution of these variables during pretraining, so theoretically it could have added the missing demographics to each persona by sampling from that joint distribution. Given that the LLM does not seem to do this on its own, it would be interesting in future work to explore this capability: adding greater variance in the generated networks by augmenting each persona with additional traits, generated by the LLM. However, one would have to be careful about exacerbating biases with this technique, as we saw with LLM-generated interests.
Second, we try providing two variables at a time: political affiliation and one of the four other variables. For each variable, its level of homophily when in a pair tends to be between its level of homophily when alone versus with all five variables (Figure B2, bottom). We also find in both of these experiments that political affiliation continues to be the dominant factor. In the single-variable experiments, political homophily is the strongest when political affiliation is provided, compared to homophily for any other variable when it is the only variable provided. In the two-variable experiments where political affiliation and one other variable is provided, political homophily is always stronger than the other variable’s homophily.
Degree per demographic group.
Figure B3: Degree distribution per demographic group, pooled over the 30 networks generated by the Sequential method.
$n$ Mean degree Gini coef. GS Index
Overall 50 0.149 (0.002) 0.347 (0.005)
Man 24 0.131 (0.003) 0.281 (0.009) 0.368 (0.011)
Woman 26 0.166 (0.004) 0.370 (0.008) 0.299 (0.009)
Under 30 14 0.147 (0.005) 0.321 (0.010) 0.524 (0.009)
30-59 19 0.150 (0.004) 0.336 (0.009) 0.543 (0.007)
60+ 17 0.150 (0.005) 0.328 (0.009) 0.541 (0.006)
Hispanic 10 0.114 (0.004) 0.259 (0.013) 0.426 (0.017)
White 33 0.160 (0.003) 0.335 (0.008) 0.330 (0.008)
Catholic 11 0.157 (0.006) 0.354 (0.013) 0.511 (0.009)
Protestant 19 0.163 (0.004) 0.352 (0.010) 0.475 (0.011)
Unreligious 20 0.132 (0.004) 0.276 (0.008) 0.487 (0.009)
Democrat 24 0.157 (0.004) 0.355 (0.009) 0.189 (0.011)
Republican 26 0.143 (0.005) 0.313 (0.007) 0.210 (0.014)
Table B3: Degree distribution and Gini-Simpson (GS) diversity index per demographic group, for all groups with at least 10 personas in our set of 50 personas. For each statistic, we compute its value per network, and report the mean value and standard error (in parentheses) over the 30 networks generated by the Sequential method.
In the main text, we discussed and visualized degree distributions over all nodes in the network (Figure 4), but degree could also differ across demographic groups. In Figure B3, we visualize each group’s degree distribution, for all groups with at least 10 personas in our set of 50 personas. As in Figure 4, the distribution consists of degrees for all nodes in the group, pooled over the 30 networks generated by the Sequential method. We see minor differences between groups: for example, in the generated networks, women have slightly higher degree on average than men (0.166 vs. 0.131, Table B3). This could be explained by two factors: first, there are slightly more women ($n$=26) than men ($n$=24) in our set of 50 personas, and second, women show higher levels of same-group preference, as discussed in the main text (same-group ratios of 1.43 vs 0.99, Figure 6). We also find that White personas have slightly higher average degree than Hispanic personas (0.160 vs 0.114), which could also be explained by there being more White personas ($n$=33) vs. Hispanic personas ($n$=10), and higher same-group ratios among White personas compared to Hispanic personas (1.20 vs. 1.02).
Beyond using the mean to describe the degree distribution, we can also use the Gini coefficient as a measure of the inequality in degree. For a set of values $x_{1},x_{2},\cdots,x_{n}$, the Gini coefficient is defined as
$G=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}|x_{i}-x_{j}|}{2n^{2}\bar{x}},$ (3)
where $\bar{x}$ is the mean of the values. First, when we compute the Gini coeffient over all nodes, we find a higher coefficient (more inequality) when using the Sequential method ($G=0.347$), compared to the Local method ($G=0.261$) or Global method ($G=0.255$). This aligns with our earlier results showing that Sequential can better capture the long-tailed degree distribution of real social networks. Per group, we find similar results to what we found for degree distribution: higher average degree tends to correspond to higher Gini coefficient, since the higher average tends to be driven by longer tails, which results in greater inequality. As described in Table B3, women compared to men, White compared to Hispanic, Protestant and Catholic compared to Unreligious, and Democrat compared to Republican all have slightly higher average degrees and Gini coefficients.
Diversity per demographic group.
We can also measure rates of homophily and diversity per group. In Figure 6, we visualized observed-to-expected ratios to capture same-group preferences (the diagonal) and cross-group preferences (the off-diagonal). This analysis revealed that not all cross-group relations are equally unlikely; for example, adjacent age groups were likelier to have cross-group relations. It also revealed that not all same-group preferences were equally strong; for example, we saw stronger preferences within women compared to men or within Catholic compared to Unreligious.
Related to homophily, we can also measure the diversity in a persona’s 1-hop neighbors (i.e., their friends). We use the Gini-Simpson (GS) index to measure diversity, which is defined in terms of a set of groups, $\mathcal{G}=\{g_{1},g_{2},\cdots,g_{k}\}$, and the proportion $p_{g}$ of values that belong to each group:
$GS=1-\sum_{g\in\mathcal{G}}p_{g}.$ (4)
For a given node $i$, we define its 1-hop diversity with respect to a demographic variable by computing the GS index over their friends’ group identities for that variable. In Table B3, we report the average 1-hop diversity for members in a group, with respect to that group’s demographic variable (e.g., the average 1-hop gender diversity for men). As expected, groups with higher diversity correspond to those with lower same-group ratios in Figure 6: men compared to women, Hispanic compared to White, Catholic compared to Protestant and Unreligious, and Republican compared to Democrat. The three age groups are similar to each other for both measures.
Individuals and intersections of group identities.
We found that LLMs overemphasize political homophily overall and that each political group, Democrats and Republicans, shows strong same-group preferences (Figure 6) and low diversity in 1-hop neighbors (Table B3). However, how does political preference and diversity vary over individuals, as their other group identities intersect with their political affiliation? To analyze this, we use individual-level metrics: the GS index to measure individual-level diversity in 1-hop neighbors and an individual’s“Democrat lean”, i.e., the proportion of the individual’s 1-hop neighbors who are Democrats, as an easily interpretable measure.
We find large within-group variability over individuals: for example, the persona with the lowest 1-hop political diversity (measured by GS index) is a woman, age 75, White, Unreligious, and Democrat (mean=0.111, SE=0.032), while the persona with the highest 1-hop political diversity is a woman, age 38, White, Protestant, and Democrat (mean=0.353, SE=0.028). Relatedly, we find that the lowest-diversity persona has a very high Democrat lean (mean=0.908, SE=0.030), while the highest-diversity persona’s lean is less extreme (mean=0.722, SE=0.028). Since both individuals are women, White, and Democrat, we can attribute their differences to their age and religion. From pure numbers, we would not expect religion to play a major role in their differences, since the share of Democrats vs. Republicans is about the same for personas that identify as Unreligious (10 vs. 10) and as Protestant (10 vs. 9). However, it is possible that the LLM expects that an Unreligious Democrat’s preference for befriending Democrats is stronger than a Protestant Democrat’s preference. Age could also partially explain their differences: among personas within 10 years of 75 (65-85), there are 6 Democrats and 3 Republicans, but among personas within 10 years of 38 (28-48), there are 8 Republicans and 5 Democrats, which could help to explain why the former persona has a stronger Democrat lean and less political diversity in friends.
These results reveal that, even though we see strong political homophily and low political diversity overall, the LLM generates some variability across individuals, due to the intersection of their political affiliation with other demographic identities, such as religion or age.
@@
Figure B4: Visualizing demographic homophily under three methods: “Seq.” (only demographic information), “Seq. w interests” (demographic information and interests), and “Seq. only interests” (only interests). The LLM places the largest emphasis on political homophily in all cases. Ratios below 1 (marked by the grey line) indicate homophily, with lower ratios indicating more homophily. We visualize mean and standard error (in black) and individual data points corresponding to each network. All model results shown here use the Sequential method and GPT-3.5-Turbo.
@@
Table B4: Top 10 interests per demographic group.
@@
In Figure B4, we visualize demographic homophily under three versions of the Sequential method: when only demographic information is provided, when demographic information and interests are provided, and when only interests are provided. In all cases, the LLM places the largest emphasis on political homophily.